OpenStudy (anonymous):

my freind wants to know what the line -1,2 and 3.-3 would be if grown or "dialated" by a factor of two.

OpenStudy (anonymous):

@e.mccormick @Kit_Kat_Nat_<3 @love10129151 @sammixboo

OpenStudy (e.mccormick):

Line or segment? A line goes on forever, so \(2\times\infty = \infty\) For everything else, the scale is just multiplied into the points.

OpenStudy (anonymous):

Thanks!